OpenStudy (compassionate):

Hi, Triangle STU is dilated to form new triangle VWX. If angle S is congruent to angle V, what other information will prove that the two triangles are similar? Side ST is congruent to side VW. Angle T is congruent to angle V. Side US is congruent to side XV. Angle U is congruent to angle X. Do not give me the answer. Thank you.

OpenStudy (kc_kennylau):

Similarity: AAA Ratios of 2 sides inc < 3 sides prop

OpenStudy (kc_kennylau):

(NB: AA can be derived from AAA using < sum of Δ)

OpenStudy (kc_kennylau):

From that we conclude: AA needs two information Ratios of 2 sides inc < needs three information 3 sides prop needs three information

OpenStudy (compassionate):

If we concluded that X is congruent to U, then we can say that this triangle is similar by AA.

OpenStudy (kc_kennylau):

yep

OpenStudy (compassionate):

Hi, Lets take a short break from that and revamp to this concern "Side AC = 4, side BC = 6, side DF = 4, and side EF = 6. What side corresponds to side AB and can be used to show that ΔABC ≅ ΔDEF by SSS? (Enter your answer using letters only)" Now, I would do AC/BC = DF/EF Then I would plug them in: 4/6 = 4/6 Cross multiply 24 = 24 However, I am confused. What does it mean "what side corresponds to aB" Wouldn't it be AC and BC that correspond to AB?

OpenStudy (compassionate):

@kc_kennylau , get over here ;D

OpenStudy (kc_kennylau):

lolz viewing multiple problems at once ain't an easy thing...

OpenStudy (compassionate):

Do mine first, and you won't regret it.

OpenStudy (kc_kennylau):



OpenStudy (kc_kennylau):

DE corresponds to AB

OpenStudy (kc_kennylau):

EF corresponds to BC

OpenStudy (kc_kennylau):

DF corresponds to AC

OpenStudy (compassionate):

I see, I see. Now, lets turn 180 degrees, and do this: >>> http://prntscr.com/2cromn (Open) It would be 5/7 = b/16 Which would get 8b = 80 Which b = 10

OpenStudy (compassionate):

5/8 = b/16 That's what I mean.

OpenStudy (kc_kennylau):

http://www.mathwarehouse.com/geometry/similar/triangles/images/sides-of-similar-triangles-are-proportional.png ok it seems that you cannot use internal ratios...

OpenStudy (kc_kennylau):

b/5=16/8

OpenStudy (compassionate):

Right. You get 8b = 80 8/80 = 10 B = 10

OpenStudy (kc_kennylau):

yep

OpenStudy (compassionate):

Hi, @kc_kennylau http://prntscr.com/2crtdz Thank you.

OpenStudy (kc_kennylau):

work out the ratios of the corresponding sides first?

OpenStudy (compassionate):

Demonstrate, please.

OpenStudy (kc_kennylau):

10/52=5/26 2.5/13=5/26 12.5/62.5=5/25

OpenStudy (kc_kennylau):

therefore EF should be 65

OpenStudy (kc_kennylau):

sorry i have to go now :/

OpenStudy (compassionate):

Bye! THANK YOU!

OpenStudy (compassionate):

@kc_kennylau , One more :) A kite is a quadrilateral with two pairs of adjacent, congruent sides. Prove the two angles between the non-congruent sides are congruent. Be sure to create and name the appropriate geometric figures. This figure does not need to be submitted.